MODS1 Commissioning Run 3: UTC 2010 December 18

Observer(s): R. Pogge and C. Henderson (OSU), M. Pedani (LBTO)
Telescope Operator: D. Gonzalez (LBTO)
Telescope Support: J. Little (LBTO)
Instrument Support: J. Morris (LBTO)

Summary

Internet and phones down most of the day because the main fiber optics link between Bonita and Tucson was cut. Service was not restored until 2UTC. Astronomers still slowly recovering from internet withdrawl...

Snow last night, 4-6 inches on the roads, lots of icy rime on everything at dawn, but today we got Sun and clouds enough of the time to clear the dome. Humidity started to drop towards evening, but still a little dicey to open up. Also the inside is 6°C warmer than outside because of the front, so we're going to be way out of equilibration for at least the first half of the night.

The night deteriorated quickly into high humidity, with icy fog blowing past later on. We took the time to do various software tests, including testing the IMCSLOCK procedure used to align the IMCS after a big move or instrument configuration change.

In addition to software testing, we took some test images and measured photon transfer curves for the blue and red CCDs to measure their gain and readout noise on-telescope.

Preparations

Dewars filled at 1600. Hold times still remain >24h

The vent line for the red camera dewar was starting to slip back inside the instrument enclosure with only the inner plastic vent line protruding and the insulating foam noodle no longer visible. RWP adjusted it so the end of the foam noodle is sticking out ~5cm beyond the end of the side-panel "stovepipe", and he afixed it with some aluminum tape as it was loose. Blue dewar vent line is fine. We should review how the vent lines are held in place at the enclosure stovepipes.

Last night's bug with the wacky exposure progress bar behavior for 0.5s exposures was identified and fixed. Implemented and tested the IMCSLOCK procedure that internalizes the "imcscen.pro" script procedure we had been using thus far to lock-on and align the IMCS. Default parameters and tuning of the IMCS were based on the Sept and November on-sky and daytime test. Will test this tonight if the weather is poor. This provides us with an efficient way to realign the IMCS inside scripts or from the command line.

Continued developing and testing procedure (.pro) and calibration (.cal) scripts for use with MODS. Have also tested meta-scripting (scripts of scripts) with good success.

Details

All times are UTC unless otherwise noted.

0225 - Humidity 77%, Temp -4°C outside, +3°C inside, wind 12m/sec from the SW, humiding rising but slower than for the past half hour. We are closed right now. The inside/outside differential is +7°C, so thermal issues are going to loom large.

0300 - Humidity 80%, Temp -4°C outside, +4°C inside, Wind 15m/s from the SW. Still not OK to open the vent doors and start equilibrating, high clouds keep coming over us.

0400 - Humidity 91%, Temp -5°C outside, +4°C inside, Wind 11m/s from SW. High clouds continue to pass overhead.

0500 - Humidity 97%, Temp -5°C outside, +4°C inside, wind 12-14m/s out of the SW. Clouds still blowing overhead.

0600 - Humidity 100%, Temp -5.5°C outside, +4°C inside, wind 5-10m/s out of the SSW. Very cool frost web formed on the VATT all-sky camera once the humidity hit 100%. Actually mostly clear to the eye in the bright moonlight but still stuff blowing in from the west.

0700 - Humidity 100%, Temp -5.5°C outside, +4°C inside, wind 5-12m/s out of the SW. Bits of cloud blowing past at ground level visible in moonlight.

0800 - Humidity 100%, Temp -5°C outside, +4°C inside, wind 8-15m/s from the S. Blowing clouds and clear, but no precipitation.

0900 - Humidity 100%, Temp -5°C outside, +4°C inside, wind 5-10m/s from S. Inside a cloud, fog is freezing to most exposed surfaces.

1000 - Clear, photometric, 0.3-arcsec seeing. No? OK, 100% Humidity, Temp -4.5°C out, +4°C in, wind 5-15m/s from the S. Sitting in freezing fog.

1100 - Humidity 100%, Temp -4.5°C outside, +4°C inside, Wind gusting to 16m/ from the SSW. In a blowing cloud with rime ice forming on most surfaces.

1200 - Humidity 100%, Temp -4.5°C out, +4°C in, wind 15-18m/s from the SW. Out of the cloud and some stars visible.

-- RichardPogge - 18 Dec 2010
Topic revision: r7 - 19 Dec 2010, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback